OpenStudy (anonymous):

what is the simplified form? (4/7)^3 a.343/64 b.21,952 c.407 d.64/343

OpenStudy (anonymous):

64 and 147

OpenStudy (anonymous):

soo. whats the answer ?

OpenStudy (anonymous):

"simplified" ??

OpenStudy (anonymous):

i guess that means " compute" right?

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=%284%2F7%29^3

OpenStudy (anonymous):

its D

OpenStudy (ahsome):

\[\left(\frac{4}{7}\right)^3\]\[=\frac{64}{343}\]Since they don't share any common factors, that is the best you can simplify. So the answer is D

OpenStudy (anonymous):

i just said that stop waisting peoples times

OpenStudy (ahsome):

Woah, calm down @prowrestler. I'm just explain why you're right.